# Cross-positive linear maps, positive polynomials and sums of squares

### v1

@@ 5. Algorithms and Examples
Each biquadratic form $f\in\mathbb{R}[{\tt x},{\tt y}]_{2,2}$ that is modulo $I$ a nonnegative polynomial but not a sum of squares yields an example of a “proper” cross-positive map $A:M_{n}(\mathbb{R})_{\textrm{sym}}\to M_{n}(\mathbb{R})_{\textrm{sym}}$, cf. Proposition 2.5. In this section we specialize the Blekherman-Smith-Velasco algorithm ([BSV16, Procedure 3.3]; see also [KMŠZ19] for a specialization in the context of positive linear maps) to produce many such examples from random input data. Throughout this section we fix $n\geq 2$.
First observe that biquadratic forms are in bijective correspondence with quadratic forms on the Segre variety (cf. [BSV16, Example 5.6]). Let $\mathbb{P}^{n-1}$ denote the complex $(n-1)$-dimensional projective space and let $\sigma_{n}:\mathbb{P}^{n-1}\times\mathbb{P}^{n-1}\to\mathbb{P}^{n^{2}-1},$ $([x_{1}:\ldots:x_{n}],[y_{1}:\ldots:y_{n}])\mapsto[x_{1}y_{1}:x_{1}y_{2}:\ldots:x_{1}y_{n}:\ldots:x_{n}y_{n}]$ be the Segre embedding. Its image $\sigma_{n}(\mathbb{P}^{n-1}\times\mathbb{P}^{n-1})=V_{\mathbb{C}}(I_{n})$ is the complex zero locus of the ideal $I_{n}\subseteq\mathbb{R}[{\tt z}]:=\mathbb{R}[{\tt z}_{11},{\tt z}_{12},\ldots,{\tt z}_{1n},\ldots,{\tt z}_{nn}]$ generated by all $2\times 2$ minors of the matrix $\left({\tt z}_{ij}\right)_{i,j}$. The complexification $I_{n}^{\mathbb{C}}\subseteq\mathbb{C}[{\tt z}]$ of the ideal $I_{n}$ is radical [Har92, p. 98] and thus consists of all polynomials vanishing on $\sigma_{n}(\mathbb{P}^{n-1}\times\mathbb{P}^{n-1})$. It is also well known that $\sigma_{n}(\mathbb{P}^{n-1}\times\mathbb{P}^{n-1})$ is smooth [Har92, p. 184-185] and of degree $\binom{2n-2}{n-1}$ [Har92, p. 233].
@@
We first show that $J_{n}^{\mathbb{C}}/I_{n}^{\mathbb{C}}$ is a radical ideal in $\mathbb{C}[{\tt z}]/I_{n}^{\mathbb{C}}$. Let $f\in\mathbb{C}[{\tt z}]/I_{n}^{\mathbb{C}}$ satisfy $f^{2}\in J_{n}^{\mathbb{C}}/I_{n}^{\mathbb{C}}$. Since $I_{n}^{\mathbb{C}}$ is radical ideal and $V_{\mathbb{C}}(I_{n})$ is the image of the Segre embedding, the Segre embedding induces an injective homomorphism betwen coordinate rings $\tilde{\sigma}_{n}^{\#}:\mathbb{C}[{\tt z}]/I_{n}^{\mathbb{C}}\to\mathbb{C}[{\tt x},{\tt y}]$ sending ${\tt z}_{ij}+I_{n}^{\mathbb{C}}\mapsto{\tt x}_{i}{\tt y}_{j}$. Clearly, $\tilde{\sigma}_{n}^{\#}(J_{n}^{\mathbb{C}}/I_{n}^{\mathbb{C}})\subseteq I^{\mathbb{C}}$, so $\left(\tilde{\sigma}_{n}^{\#}(f)\right)^{2}=\tilde{\sigma}_{n}^{\#}(f^{2})\in I^{\mathbb{C}}$. Since $I^{\mathbb{C}}$ is a prime ideal in $\mathbb{C}[{\tt x},{\tt y}]$, it follows that $\tilde{\sigma}_{n}^{\#}(f)\in I^{\mathbb{C}}$. Let
@@
Finally, $\mathbb{C}[{\tt z}]/I_{n}^{\mathbb{C}}$ is a domain, and $J_{n}^{\mathbb{C}}/I_{n}^{\mathbb{C}}$ is a radical ideal. It follows from this that $\left(\mathbb{C}[{\tt z}]/I_{n}^{\mathbb{C}}\right)/\left(J_{n}^{\mathbb{C}}/I_{n}^{\mathbb{C}}\right)\cong\mathbb{C}[{\tt z}]/J_{n}^{\mathbb{C}}$ is reduced (without nilpotents), hence $J_{n}^{\mathbb{C}}$ is radical ideal in $\mathbb{C}[{\tt z}]$. ∎
@@
Recall from Lemma 2.3 that a biform $f\in\mathbb{R}[{\tt x},{\tt y}]_{2,2}$ is a sum of squares modulo $I$ iff it is a sum of squares of biforms from $\mathbb{R}[{\tt x},{\tt y}]_{1,1}$ modulo $I$.
@@
To prove the implication ()($\Rightarrow$)( ⇒ ) let
@@
It remains to prove the implication ()($\Leftarrow$)( ⇐ ). Since $f$ is in the image of $\sigma_{n}^{\#}$, it follows from
@@
where $j_{1}\in\mathbb{N}$ and $[h_{i}]$ is the equivalence class of $h_{j}\in\mathbb{R}[{\tt z}]$ in $\mathbb{R}[{\tt z}]/J_{n}$, that
@@
is a sum of squares in $\mathbb{R}[{\tt x},{\tt y}]/I$ which proves ()($\Leftarrow$)( ⇐ ). ∎
@@
Next, we compute the Hilbert polynomial for $V_{\mathbb{C}}(J_{n})$. We follow the proof of the analogous result for the Segre variety in [Har92, p. 234]. The space of polynomials of degree $d$ in $\mathbb{C}[{\tt z}]/J_{n}^{\mathbb{C}}$ is isomorphic, via the restiction of the homomorphism $\sigma_{n}^{\#}$, to the space of polynomials of bidegree $(d,d)$ in $\mathbb{C}[{\tt x},{\tt y}]/I^{\mathbb{C}}$. Its dimension is therefore
(${n+d-1\choose n-1}^{2}-{n+d-2\choose n-1}^{2}=\left(\frac{(d+1)\cdots(n+d-2)}{(n-1)!}\right)^{2}(n-1)(n+2d-1).$
@@
It remains to prove smoothness. Note that the group $\operatorname{GL}_{n}$ acts on the variety $V_{\mathbb{C}}(J_{n})$ of rank 1 traceless matrices by conjugation. Using the Jordan normal form we see that the action is transitive, so it suffices to prove that $e_{1}e_{2}^{T}$ is a smooth point of $V_{\mathbb{C}}(J_{n})$. To show this we use the Jacobian criterion. Let $\operatorname{Jac}(e_{1}e_{2}^{T})$ be the Jacobian matrix for $V_{\mathbb{C}}(J_{n})$ at $e_{1}e_{2}^{T}$. The generators of the ideal $J_{n}^{\mathbb{C}}$ are ${\tt z}_{ij}{\tt z}_{kl}-{\tt z}_{il}{\tt z}_{kj}$ where $i\neq k$ and $j\neq l$, and $\sum_{i=1}^{n}{\tt z}_{ii}$. The gradient of ${\tt z}_{ij}{\tt z}_{kl}-{\tt z}_{il}{\tt z}_{kj}$ in $e_{1}e_{2}^{T}$ is zero if $i\neq 1$ and $k\neq 1$ or if $j\neq 2$ and $l\neq 2$. On the other hand, the gradient of ${\tt z}_{12}{\tt z}_{kl}-{\tt z}_{1l}{\tt z}_{k2}$ in $e_{1}e_{2}^{T}$ is $e_{k}e_{l}^{T}$ for $k\neq 1$ and $l\neq 2$, and the gradient of $\sum_{i=1}^{n}{\tt z}_{ii}$ is $\sum_{i=1}^{n}e_{i}e_{i}^{T}$. Clearly, $\sum_{i=1}^{n}e_{i}e_{i}^{T}\in\mathbb{C}^{n^{2}}$ is not a linear combination of $e_{k}e_{l}^{T}$ with $k\neq 1$ and $l\neq 2$, so $\operatorname{rank}\operatorname{Jac}(e_{1}e_{2}^{T})=(n-1)^{2}+1$ and $\dim\ker\operatorname{Jac}(e_{1}e_{2}^{T})=2n-2$. It follows that the (projective) tangent space to $V_{\mathbb{C}}(J_{n})$ at $e_{1}e_{2}^{T}$ is $(2n-3)$-dimensional, which shows that $e_{1}e_{2}^{T}$ is smooth. ∎
@@
For $n\geq 3$, [BSV16, Procedure 3.3] yields an explicit construction of nonnegative quadratic forms from $\mathbb{R}[{\tt z}]/{J_{n}}$ that are not sums of squares forms starting from random input data. We now present this procedure specialized to our context.
@@ 1. Step 1
      (Note that we have omitted $z^{(e+1)}$.) The corresponding linear form $h_{0}$ is
@@ 2. Step 2
      (The latter condition ensures $v$ is a symmetric tensor in $\mathbb{R}^{n^{2}}\otimes\mathbb{R}^{n^{2}}$. Note also that we have omitted the point $z^{(e+1)}$.)
@@
      $\operatorname{span}$ ($\Big{(}\left\{v_{i}\otimes v_{j}+v_{j}\otimes v_{i}\colon 0\leq i\leq j\leq d\right\}$
@@
      then repeat Step 2Step 2.2. Otherwise the quadratic form
@@ 3. Step 3
The main ingredient in the proof is the theory of minimal degree varieties as developed in [BSV16]. Since $V_{\mathbb{C}}(J_{n})$ is not of minimal degree for $n\geq 3$ by Proposition 5.3, $\operatorname{Sos}(V(J_{n}))\subsetneq\operatorname{Pos}(V(J_{n}))$. Hence results of [BSV16, Section 3] apply; their Procedure 3.3 adapted to our set-up is Algorithm 5.5. While Step 1 and Step 3 follow immediately from the corresponding steps in [BSV16, Procedure 3.3], we note for Step 2 that “vanishing to the second order at $z^{(i)}$” means $f(z^{(i)})=0$ and $\nabla f(z^{(i)})\in\operatorname{span}\left\{\nabla g_{j}(z^{(i)})\colon 0\leq j\leq\binom{n}{2}^{2}\right\}$. Since $f\not\in\mathfrak{a}^{2}$, the quadratic form $\delta f+\sum_{i=0}^{d}h_{i}^{2}$ is never a sum of squares, while it is nonnegative on $V(J_{n})$ for sufficiently small $\delta>0$ by the positive definiteness of the Hessian of $\sum_{i=0}^{d}h_{i}^{2}$ at its real zeros $z^{(1)},\ldots,z^{(e)}$, see the proof of the correctness of Procedure 3.3 in [BSV16]. We note that the verification in Step 1Step 1.2 is computationally difficult, but since all steps in the algorithm are performed with random data, all the generic conditions from [BSV16, Procedure 3.3] are satisfied with probability 1. Hence, Algorithm 5.5 works well with probability 1 without implementing verifications.
5.2. Implementation
Step 1 and Step 2 are easily implemented as they only require linear algebra. (The verification in Step 1Step 1.2 could be performed using Gröbner basis if $n$ is small, but is with high probability satisfied on random input data.) On the other hand, Step 3 is computationally difficult; testing nonnegativity even of low degree polynomials is NP-hard, cf. [LNQY10].
Several algorithms are available to check nonnegativity of polynomials. Those using symbolic methods such as quantifier elimination or cylindrical algebraic decomposition only work for small problem sizes. We employ numerical methods based on the Positivstellensatz [BCR98]. To reduce the number of equality constraints, we rewrite the quadratic form $F_{\delta}({\tt z})$ from (5.3) into ${\tt x},{\tt y}$ variables, obtaining a biquadratic form we denote by a slight abuse of notation by $F_{\delta}({\tt x},{\tt y})$.
@@ 5.2. Implementation
Assume (ii) holds. From (5.4) it follows that $f\geq 0$ on $S=V(I)\setminus V(\sigma_{1})$. Since $V(I)$ is irreducible, $S$ is Zariski dense in $V(I)$. Since $S$ is also open in $V(I)$, it is dense in $V(I)$ also in the Euclidean topology and hence (i) holds. Conversely, suppose (i) holds. By the Positivstellensatz (e.g. [BCR98, Corollary 4.4.3]), there is $m\in\mathbb{N}$ and sums of squares $\sigma_{1},\sigma_{2}$ such that $f\sigma_{1}-\sigma_{2}-f^{2m}\in I$. Assume $\sigma_{1}=0$ on $V(I)$, then $\sigma_{1}\in I$ since $I$ is the vanishing ideal of $V(I)$ (see §2), whence $\sigma_{2}+f^{2m}\in I$. Thus, again by the real radical property of $I$, $f\in I$. In this case we may simply pick $\sigma_{1}=1$ and $\sigma_{2}=0$. ∎
We apply Proposition 5.6 to $F_{\delta}$ from (5.3) to search for a $\delta>0$ making $F_{\delta}({\tt x},{\tt y})$ nonnegative on $V(I)$. Let $\delta>0$ be fixed and suppose the degree of $\sigma_{1}$ is $\leq 2d$. Then the ideal constraint in (5.4) immediately converts into a linear matrix inequality and thus feeds into a semidefinite program (SDP) that can be solved with standard solvers [WSV00]. (Here homogeneity of $F_{\delta}$ and $I$ enter. Both $\sigma_{j}$ can be assumed to be homogeneous, and $\deg\sigma_{2}\leq 2d+4$.) To implement the non-equality constraint in (5.4), we pick a random point $(x_{0},y_{0})\in V(I)$ and set $\sigma_{1}(x_{0},y_{0})=1$. Our implementation uses bisection, sets $d=1$ and starts with, say, $\delta=1$. Then solve the described feasibility SDP. If it has a solution, stop. If not, replace $\delta$ by $\delta/2$ and try again. If no solution has been found with $\delta$ greater than some prescribed tolerance, increase $d$, and reset $\delta=1$. Then repeat the process. By Proposition 5.6 and the construction of $F_{\delta}$ the algorithm will eventually produce a certificate of nonnegativity for some $\delta>0$. We refer to [Bha23] for a numerical comparison of polynomial optimization choices for a similar problem.
As in [KMŠZ19] (see also [Bha23]) it is often possible to apply standard techniques [PP08, CKP15] to rationalize obtained sum of squares certificates.

### v2

@@ 5. Algorithms and Examples
Each biquadratic form $f\in\mathbb{R}[{\tt x},{\tt y}]_{2,2}$ that is modulo $I$ a nonnegative polynomial but not a sum of squares yields an example of a “proper” cross-positive map $A:M_{n}(\mathbb{R})\to M_{n}(\mathbb{R})$, cf. Proposition 2.6. In this section we specialize the Blekherman-Smith-Velasco procedure ([6, Procedure 3.3]; see also [22] for a specialization in the context of positive linear maps) to produce many such examples from random input data. Throughout this section we fix $n\geq 2$.
Observe that biquadratic forms are in bijective correspondence with quadratic forms on the Segre variety (cf. [6, Example 5.6] or [41, Lemma 3.13]). Let $\mathbb{P}^{n-1}$ denote the complex (n1)(n-1)-dimensional projective space and let $\sigma_{n}:\mathbb{P}^{n-1}\times\mathbb{P}^{n-1}\to\mathbb{P}^{n^{2}-1},$ $([x_{1}:\ldots:x_{n}],[y_{1}:\ldots:y_{n}])\mapsto[x_{1}y_{1}:x_{1}y_{2}:\ldots:x_{1}y_{n}:\ldots:x_{n}y_{n}]$ be the Segre embedding. Its image $\sigma_{n}(\mathbb{P}^{n-1}\times\mathbb{P}^{n-1})=V_{\mathbb{C}}(I_{n})$ is the complex zero locus of the ideal $I_{n}\subseteq\mathbb{R}[{\tt z}]:=\mathbb{R}[{\tt z}_{11},{\tt z}_{12},\ldots,{\tt z}_{1n},\ldots,{\tt z}_{nn}]$ generated by all $2\times 2$ minors of the matrix $\left({\tt z}_{ij}\right)_{i,j}$. The complexification $I_{n}^{\mathbb{C}}\subseteq\mathbb{C}[{\tt z}]$ of the ideal $I_{n}$ is radical [18, p. 98] and thus consists of all polynomials vanishing on $\sigma_{n}(\mathbb{P}^{n-1}\times\mathbb{P}^{n-1})$. It is also well known that $\sigma_{n}(\mathbb{P}^{n-1}\times\mathbb{P}^{n-1})$ is smooth [18, p. 184-185] and of degree $\binom{2n-2}{n-1}$ [18, p. 233].
@@
We first show that $J_{n}^{\mathbb{C}}/I_{n}^{\mathbb{C}}$ is a radical ideal in $\mathbb{C}[{\tt z}]/I_{n}^{\mathbb{C}}$. Let $f\in\mathbb{C}[{\tt z}]/I_{n}^{\mathbb{C}}$ satisfy $f^{2}\in J_{n}^{\mathbb{C}}/I_{n}^{\mathbb{C}}$. Since $I_{n}^{\mathbb{C}}$ is radical ideal and $V_{\mathbb{C}}(I_{n})$ is the image of the Segre embedding, the Segre embedding induces an injective homomorphism between coordinate rings $\tilde{\sigma}_{n}^{\#}:\mathbb{C}[{\tt z}]/I_{n}^{\mathbb{C}}\to\mathbb{C}[{\tt x},{\tt y}]$ sending ${\tt z}_{ij}+I_{n}^{\mathbb{C}}\mapsto{\tt x}_{i}{\tt y}_{j}$. Clearly, $\tilde{\sigma}_{n}^{\#}(J_{n}^{\mathbb{C}}/I_{n}^{\mathbb{C}})\subseteq I^{\mathbb{C}}$, so $\left(\tilde{\sigma}_{n}^{\#}(f)\right)^{2}=\tilde{\sigma}_{n}^{\#}(f^{2})\in I^{\mathbb{C}}$. Since $I^{\mathbb{C}}$ is a prime ideal in $\mathbb{C}[{\tt x},{\tt y}]$, it follows that $\tilde{\sigma}_{n}^{\#}(f)\in I^{\mathbb{C}}$. Let
@@
Finally, $\mathbb{C}[{\tt z}]/I_{n}^{\mathbb{C}}$ is a domain, and $J_{n}^{\mathbb{C}}/I_{n}^{\mathbb{C}}$ is a radical ideal. It follows that
$\left(\mathbb{C}[{\tt z}]/I_{n}^{\mathbb{C}}\right)/\left(J_{n}^{\mathbb{C}}/I_{n}^{\mathbb{C}}\right)\cong\mathbb{C}[{\tt z}]/J_{n}^{\mathbb{C}}$
is reduced (without nilpotents), hence $J_{n}^{\mathbb{C}}$ is a radical ideal in $\mathbb{C}[{\tt z}]$. ∎
@@
Recall from Lemma 2.4 that a biform $f\in\mathbb{R}[{\tt x},{\tt y}]_{2,2}$ is a sum of squares modulo $I$ if and only if it is a sum of squares of biforms from $\mathbb{R}[{\tt x},{\tt y}]_{1,1}$ modulo $I$.
@@
To prove the implication $(\Rightarrow)$ let
@@
It remains to prove the implication $(\Leftarrow)$. Since $f$ is in the image of $\sigma_{n}^{\#}$, it follows from
@@
where $j_{1}\in\mathbb{N}$ and $h_{i}]$ is the equivalence class of $h_{j}\in\mathbb{R}[{\tt z}]$ in $\mathbb{R}[{\tt z}]/J_{n}$, that
@@
is a sum of squares in $\mathbb{R}[{\tt x},{\tt y}]/I$ which proves $(\Leftarrow)$. ∎
@@
Next, we compute the Hilbert polynomial for $V_{\mathbb{C}}(J_{n})$. We follow the proof of the analogous result for the Segre variety in [18, p. 234]. The space of polynomials of degree $d$ in $\mathbb{C}[{\tt z}]/J_{n}^{\mathbb{C}}$ is isomorphic, via the restriction of the homomorphism $\sigma_{n}^{\#}$, to the space of polynomials of bidegree $(d,d)$ in $\mathbb{C}[{\tt x},{\tt y}]/I^{\mathbb{C}}$. Its dimension is therefore
${n+d-1\choose n-1}^{2}-{n+d-2\choose n-1}^{2}=\left(\frac{(d+1)\cdots(n+d-2)}{(n-1)!}\right)^{2}(n-1)(n+2d-1).$
@@
It remains to prove smoothness. Note that the group $\operatorname{GL}_{n}$ acts on the variety $V_{\mathbb{C}}(J_{n})$ of rank 1 traceless matrices by conjugation. Using the Jordan normal form we see that the action is transitive, so it suffices to prove that $e_{1}e_{2}^{T}$ is a smooth point of $V_{\mathbb{C}}(J_{n})$. To show this we use the Jacobian criterion. Let $\operatorname{Jac}(e_{1}e_{2}^{T})$ be the Jacobian matrix for $V_{\mathbb{C}}(J_{n})$ at $e_{1}e_{2}^{T}$. The generators of the ideal $J_{n}^{\mathbb{C}}$ are ${\tt z}_{ij}{\tt z}_{kl}-{\tt z}_{il}{\tt z}_{kj}$ where $i\neq k$ and $j\neq l$, and $\sum_{i=1}^{n}{\tt z}_{ii}$. The gradient of ${\tt z}_{ij}{\tt z}_{kl}-{\tt z}_{il}{\tt z}_{kj}$ in $e_{1}e_{2}^{T}$ is zero if $i\neq 1$ and $k\neq 1$ or if $j\neq 2$ and $l\neq 2$. On the other hand, the gradient of ${\tt z}_{12}{\tt z}_{kl}-{\tt z}_{1l}{\tt z}_{k2}$ in $e_{1}e_{2}^{T}$ is $e_{k}e_{l}^{T}$ for $k\neq 1$ and $l\neq 2$, and the gradient of $\sum_{i=1}^{n}{\tt z}_{ii}$ is $\sum_{i=1}^{n}e_{i}e_{i}^{T}$. Clearly, $\sum_{i=1}^{n}e_{i}e_{i}^{T}\in\mathbb{C}^{n^{2}}$ is not a linear combination of $e_{k}e_{l}^{T}$ with $k\neq 1$ and $l\neq 2$, so $\operatorname{rank}\operatorname{Jac}(e_{1}e_{2}^{T})=(n-1)^{2}+1$ and $\dim\ker\operatorname{Jac}(e_{1}e_{2}^{T})=2n-2$. It follows that the (projective) tangent space to $V_{\mathbb{C}}(J_{n})$ at $e_{1}e_{2}^{T}$ is (2n3)(2n-3)-dimensional, which shows that $e_{1}e_{2}^{T}$ is smooth. ∎
@@
For $n\geq 3$, [6, Procedure 3.3] yields an explicit construction of nonnegative quadratic forms from $\mathbb{R}[{\tt z}]/{J_{n}}$ that are not sums of squares forms starting from random input data. We now turn this procedure, specialized to our context, into a probabilistic (Las Vegas) algorithm.
@@ 1. Step 1
      (Note $z^{(e+1)}$ is omitted.) The corresponding linear form $h_{0}$ is
@@ 2. Step 2
      (The latter condition ensures $v$ is a symmetric tensor in $\mathbb{R}^{n^{2}}\otimes\mathbb{R}^{n^{2}}$. Note also the point $z^{(e+1)}$ is omitted.)
@@
      $\operatorname{span}$ $\Big{(}\left\{v_{i}\otimes v_{j}+v_{j}\otimes v_{i}\colon 0\leq i\leq j\leq d\right\}$
@@
      then repeat Step 2Step $2.$ Otherwise the quadratic form
@@ 3. Step 3
The main ingredient in the proof is the theory of minimal degree varieties as developed in [6]. Since $V_{\mathbb{C}}(J_{n})$ is not of minimal degree for $n\geq 3$ by Proposition 5.3, $\operatorname{Sos}(V(J_{n}))\subsetneq\operatorname{Pos}(V(J_{n}))$. Hence results of [6, Section 3] apply; their Procedure 3.3 adapted to our set-up is Algorithm 5.5. While Step 1 and Step 3 follow immediately from the corresponding steps in [6, Procedure 3.3], we note for Step 2 that “vanishing to the second order at $z^{(i)}$” means $f(z^{(i)})=0$ and $\nabla f(z^{(i)})\in\operatorname{span}\left\{\nabla g_{j}(z^{(i)})\colon 0\leq j\leq\binom{n}{2}^{2}\right\}$. Since $f\not\in\mathfrak{a}^{2}$, the quadratic form $\delta f+\sum_{i=0}^{d}h_{i}^{2}$ is never a sum of squares, while it is nonnegative on $V(J_{n})$ for sufficiently small $\delta>0$ by the positive definiteness of the Hessian of $\sum_{i=0}^{d}h_{i}^{2}$ at its real zeros $z^{(1)},\ldots,z^{(e)}$, see the proof of the correctness of Procedure 3.3 in [6].
5.2. Towards an implementation
Note that the verification in Step 1Step 1.2 is computationally difficult (but can be performed using Gröbner basis if $n$ is small). However, since all steps in the algorithm are performed with random data, all the generic conditions from [6, Procedure 3.3] are satisfied with probability 1. Hence, by omitting the verification in Step 1Step 1.2, Algorithm 5.5 becomes a Monte Carlo algorithm and yields a correct output with probability 1. Step 1 and Step 2 are easily implemented as they only require linear algebra. On the other hand, Step 3 is computationally difficult; testing nonnegativity even of low degree polynomials is NP-hard, cf. [26].
Several algorithms are available to check nonnegativity of polynomials. Those using symbolic methods such as quantifier elimination or cylindrical algebraic decomposition only work for small problem sizes. We employ numerical methods based on the Positivstellensatz [8]. To reduce the number of equality constraints, we rewrite the quadratic form $F_{\delta}({\tt z})$ from (5.3) into ${\tt x},{\tt y}$ variables, obtaining a biquadratic form we denote by a slight abuse of notation by $F_{\delta}({\tt x},{\tt y})$.
@@ 5.2. Towards an implementation
Assume (ii) holds. From (5.4) it follows that $f\geq 0$ on $S=V(I)\setminus V(\sigma_{1})$. Since $V(I)$ is irreducible, $S$ is Zariski dense in $V(I)$. Since $S$ is also open in $V(I)$, it is dense in $V(I)$ also in the Euclidean topology and hence (i) holds. Conversely, suppose (i) holds. By the Positivstellensatz (e.g. [8, Corollary 4.4.3]), there is $m\in\mathbb{N}$ and sums of squares $\sigma_{1},\sigma_{2}$ such that $f\sigma_{1}-\sigma_{2}-f^{2m}\in I$. Assume $\sigma_{1}=0$ on $V(I)$, then $\sigma_{1}\in I$ since $I$ is the vanishing ideal of $V(I)$ (see §2), whence $\sigma_{2}+f^{2m}\in I$. Thus, again by the real radical property of $I$, $f\in I$. In this case we may simply pick $\sigma_{1}=1$ and $\sigma_{2}=0$. ∎
We apply Proposition 5.6 to $F_{\delta}$ from (5.3) to search for a $\delta>0$ making $F_{\delta}({\tt x},{\tt y})$ nonnegative on $V(I)$. Let $\delta>0$ be fixed and suppose the degree of $\sigma_{1}$ is $\leq 2d$. Then the ideal constraint in (5.4) immediately converts into a linear matrix inequality and thus feeds into a semidefinite program (SDP) that can be solved with standard solvers [43]. (Here homogeneity of $F_{\delta}$ and $I$ enter. Both $\sigma_{j}$ can be assumed to be homogeneous, and $\deg\sigma_{2}\leq 2d+4$.) To implement the non-equality constraint in (5.4), we pick a random point $(x_{0},y_{0})\in V(I)$ and set $\sigma_{1}(x_{0},y_{0})=1$. Our implementation uses bisection, sets $d=1$ and starts with, say, $\delta=1$. Then solve the described feasibility SDP. If it has a solution, stop. If not, replace $\delta$ by $\delta/2$ and try again. If no solution has been found with $\delta$ greater than some prescribed tolerance, increase $d$, and reset $\delta=1$. Then repeat the process. By Proposition 5.6 and the construction of $F_{\delta}$ the algorithm will eventually produce a certificate of nonnegativity for some $\delta>0$. We refer to [4] for a numerical comparison of polynomial optimization choices for a similar problem.
As in [22] (see also [4]) it might be possible to apply standard techniques [32, 9] to turn obtained numerical sum of squares certificates into symbolic ones.